# Curvature-Constrained Vector Field for Motion Planning of Nonholonomic Robots

### v1

@@ Abstract
Vector fields are advantageous in handling nonholonomic motion planning as they provide reference orientation for robots. However, additionally incorporating curvature constraints becomes challenging, due to the interconnection between the design of the curvature-bounded vector field and the tracking controller under underactuation. In this paper, we present a novel framework to co-develop the vector field and the control laws, guiding the nonholonomic robot to the target configuration with curvature-bounded trajectory. First, we formulate the problem by introducing the target positive limit set, which allows the robot to converge to or pass through the target configuration, depending on different dynamics and tasks. Next, we construct a curvature-constrained vector field (CVF) via blending and distributing basic flow fields in workspace and propose the saturated control laws with a dynamic gain, under which the tracking error’s magnitude decreases even when saturation occurs. Under the control laws, kinematically constrained nonholonomic robots are guaranteed to track the reference CVF and converge to the target positive limit set with bounded trajectory curvature. Numerical simulations show that the proposed CVF method outperforms other vector-field-based algorithms. Experiments on Ackermann UGVs and semi-physical fixed-wing UAVs demonstrate that the method can be effectively implemented in real-world scenarios.  [^1] To facilitate reproducibility and further research, the source code for implementation in this paper is available at https://github.com/Y1kee/CVF-for-Nonholonomic-Motion-Planning.
Index Terms:
Motion planning, mobile robot, vector field, curvature constraint, nonholonomic constraint, saturated control.
@@ I Introduction
Motion planning, a fundamental problem in robotics, is critical to fulfill various robotic tasks, such as autonomous driving, aerial surveillance, and underwater exploration. A key task in motion planning is to provide feasible reference trajectories and control inputs to guide the robot to the target configuration. To be feasible and efficient, motion planning algorithms should take into account the robot’s kinematic constraints, among which the nonholonomic and curvature constraints are widely encountered. Nonholonomic constraints are usually imposed by the lateral velocity of robots being zero [1, 2, 3], under which the longitudinal linear velocity together with the angular velocity are regarded as the control inputs in motion planning [4, 5, 6]. Besides, as a result of the limited actuation of nonholonomic robots, such as unmanned ground vehicles (UGVs) with minimum turning radius [7, 8], unmanned aerial vehicles (UAVs) with maximum lateral acceleration [9, 10], and autonomous underwater vehicles (AUVs) with limited control moment [11, 12], the trajectory curvature of these robots is subject to an upper bound, known as the curvature constraint. The co-existence of the nonholonomic and curvature constraints poses great challenges for motion planning, due to the tightly-coupled position and orientation kinematics as well as the robots’ limited actuation.
@@
Existing motion planning methods addressing the above constraints fall into several categories, each of which has demerits hindering its practical implementation. Search-based methods are widely employed, which nevertheless require post-processing and high computational cost to handle kinematic constraints [13, 14, 15]. Optimization-based methods are capable of dealing with different constraints but often suffer from issues stemming from local minima and heavy computational burden [16, 17, 18]. Another well-studied trajectory planning algorithm is the Dubins-curves-based method, which has merits in finding curvature-bounded trajectories for nonholonomic robots; yet it only focuses on the path generation and does not involve designing control inputs for path following [19, 20, 21]. Moreover, the aforementioned methods are open-loop and hence require re-planning when deviations occur in the motion planning process.
@@
The vector field (VF)-based approaches are naturally advantageous in addressing the nonholonomic motion constraint, because a desirable VF specifies the reference orientation of the nonholonomic robot by assigning a vector to each point in the workspace. One way to generate the VF for nonholonomic robot navigation is the artificial potential field (APF) [22, 23, 24]. However, APF method involves calculating the gradient of potential functions and often suffers from local minima, probably leading to deadlock in robot’s motions. To deal with this problem, Lindemann et al. utilize cell decomposition approach to design a non-gradient VF for each polygonal cell and achieve convergence to the target position [25, 26, 27]. Inspired by the dipole-like VF, Panagou et al. derive feedback control laws to stabilize a nonholonomic robot [28], and coordinate the motion of multiple unicycles [29]. He et al. [30] further generalize the dipole-like VF to 3D motion planning and propose tracking control laws, guiding multiple nonholonomic robots to their target positions with specific heading direction. To address the issues of potential singularity commonly encountered in VFs, singularity-free VFs and control laws for fixed-wing UAVs are proposed to achieve path-following and motion coordination in [31, 32].
As to the motion planning in nonholonomic robots with bounded curvature trajectory, few existing works attempt to incorporate the curvature constraint into the VF-based motion planning. The challenges lie in at least two respects. On one hand, the construction of feasible VFs that meets the desirable curvature constraints is challenging. Given the requirement of velocity continuity in mechanics, the curvature of VFs should be not only bounded but also continuous as mentioned in [33]. Furthermore, the integral curves of the VF are expected to guide a nonholonomic robot to the target configuration in the sense of arriving at the desired position with a specified orientation. However, it is challenging to design a feasible VF such that it aligns with the target orientation when nonholonomic robots are controlled to approach the target position, since the curvature constraint limits the changing rate of orientation. For instance, the VFs proposed in [34] would converge to the desired position but fail to guarantee alignment with a specified orientation, which further highlights the difficulty of VF design in simultaneous position and orientation planning.
On the other hand, technical challenges arise when deriving control laws to track the VF under the limitations on control actuation. When the robot’s orientation is the same as the reference orientation specified by the VF, the magnitude of the VF can serve as the linear velocity of the robot while the angular velocity is determined by the changing rate of VF’s orientation[35]. However, in practice where the robot’s orientation deviates from the VF, it is difficult to track the VF while satisfying the curvature constraint, since the curvature constraint not only imposes control input bounds but also couples the linear and angular velocities. Existing works such as [36, 37] propose VFs satisfying the curvature constraint, but require overly large control inputs that possibly violate the curvature constraint. To ensure the existence of feasible control laws that track the VF under the curvature constraint, the co-design of the VF and control laws is crucial. Due to the aforementioned challenges, it remains an open problem to derive a feedback motion planning method that guides a nonholonomic robot under curvature constraint to the target configuration.
@@
In this paper, we formulate the simultaneous position and orientation planning as guiding the curvature-constrained nonholonomic robot to the target positive limit set—a positive limit set containing the target configuration. To achieve this objective, we aim to find the reference trajectories and control inputs that satisfy curvature constraint while converging to the target positive limit set. This formulation is more general and applies to real-world scenarios, including Ackermann UGV converging to a configuration, and fixed-wing UAV passing through a configuration with non-zero velocity. To address the challenges mentioned above, we propose a unified VF-based motion planning framework, under which the VF provides the curvature-bounded reference trajectories and the control laws ensure that the nonholonomic robot tracks the proposed VF under the curvature constraint.
The theoretical contributions of this paper are outlined as follows:
@@
  • We construct a curvature-constrained vector field (CVF) with a stable limit cycle as the target positive limit set. A novel VF is proposed by partitioning the workspace and embedding the blended source, sink and vortex flow fields into different regions with conveniently tunable parameters. Moreover, we derive sufficient conditions ensuring that the CVF’s integral curves have a prescribed curvature bound and serve as the reference trajectories converging to the limit cycle except at an isolated singular point. Compared with the dipole-like VF [28, 30] with singular target position and curvature-constrained VF [37, 36], this construction strategy reduces the effect of singularity during trajectory convergence, while allows for a flexible selection of CVF parameters, facilitating the co-design of CVF and control laws.
@@
  • We design novel control laws comprising the linear velocity with specified bounds adaptive to the planning objective and the saturated angular velocity, which ensures that the actual trajectory curvature is also bounded when tracking the CVF. To guarantee that the closed-loop system tracks the CVF under saturation of control inputs, we propose a state-dependent dynamic gain that responds to changes in the CVF. In this way, the orientation error between the robot and CVF is monotonically stabilized even when saturation occurs.
@@
To validate the theoretical results, we conduct and open source the implementations including the numerical simulations and hardware experiments [^1] Implementation detail and code are available at https://github.com/Y1kee/CVF-for-Nonholonomic-Motion-Planning, as summarized below.
  • We carry out numerical simulations and Monte Carlo experiments that compare the proposed method to the VF-based algorithms in [29, 35, 36, 31]. The results confirm the theoretical guarantees of our approach, achieving a 100% success rate in ensuring convergence under curvature constraints.
  • We implement the proposed planning algorithm on the Ackermann UGV featuring a minimal turning radius. The experiment results show that they successfully converge to different destinations from various initial conditions, demonstrating that the proposed feedback motion planning algorithm is robust against the measurement noise and dynamics uncertainties.
  • We extend the proposed method to 3D fixed-wing UAVs with coordinated turn constraint by deriving the attitude and thrust control laws. In the hardware-in-the-loop (HIL) experiments with target configuration switching during the flight, the UAV reaches its destination under the proposed control laws. The HIL results highlight that our algorithm given by closed-form expressions enables real-time motion planning despite limited onboard computational resources.
The remainder of this paper is structured as follows. Section II presents the nonholonomic kinematics with curvature constraint and formulates the motion planning problem. Section III constructs the CVF and presents its desired properties of convergence and bounded curvature. Section IV proposes the saturated control laws with dynamic gain and shows the almost global convergence under curvature constraint. Section V presents numerical simulations to verify the theoretical results and conduct comparative study with other existing VF-based motion planners. Section VI further examines the effectiveness of the proposed method via experiments with the Ackermann UGV and semi-physical fixed-wing UAV for robots with curvature-bounded trajectories. Finally, Section VII concludes the paper.
@@ II-A Modeling of Kinematics and Constraints
In this paper, we let $\bm{g}=(\bm{p},\theta)$ denote the configuration of a 2D mobile robot, where $\bm{p}=[x,y]^{\intercal}\in\mathbb{R}^{2}$ is the position and $\theta\in[0,2\pi)$ is the orientation. For the considered configuration, the nonholonomic constraint is imposed by confining the translation motion along the heading direction of the $x$-axis of the body-fixed frame. In this case, the most commonly-used nonholonomic kinematics are written as
@@
where $v_{x}$ is the linear velocity along the $x$-axis of the body-fixed frame and $\omega$ is the angular velocity. Moreover, $v_{x}$ and $\omega$ are considered as the control inputs to be determined by the motion planner.
In order to avoid steep manipulation or sharp turn in robot control, curvature constraint should be considered to impose an upper bound on the trajectory curvature. The curvature constraint for the nonholonomic kinematics (1) can be depicted by
@@
where $\kappa$ is the curvature of robot’s trajectory and $\bar{\kappa}$ is the maximum curvature. Additionally, the minimum turning radius of the nonholonomic robot is denoted by
@@
In the following sections, we propose a motion planning algorithm and conduct rigorous analysis based on the aforementioned kinematic model. This model captures the essential constraints and serves as a foundation extendable to more complex robotic systems, such as UGVs with Ackermann steering geometry and fixed-wing UAVs with banking turn control featuring cruising speed constraints in our experiments.
@@ II-B Problem Formulation
In real-world scenarios, space robots are usually required to arrive at the target position with predefined orientation to accomplish specific tasks. Under circumstances where robots are imposed by a positive lower bound on the velocity due to the mechanical constraints or task requirements, the convergence to a certain configuration is not always feasible. In other words, these robots are required to pass by rather than converging to a specified location. For a rigorous problem formulation, we employ the concept of positive limit set and define the target positive limit set as follows.
@@
According to the above definitions, if the controlled robot converges to the target positive limit set, then the robot will be able to arrive at the target configuration.
The objective considered in this paper includes the following: for a given target configuration $\bm{g}_{d}=(\bm{p}_{d},\theta_{d})$, we aim to propose a motion planning algorithm and construct curvature-bounded reference trajectories. Then, we will propose tracking control laws involving angular velocity $\omega$ and linear velocity $v_{x}$ such that the curvature constrained nonholonomic system (1)(2) converges to the target positive limit set $\mathcal{L}_{d}^{+}(\bm{g}_{0})\ni\bm{g}_{d}$, i.e.,
@@
The motivation of problem formulation can be explained twofold. For nonholonomic robots like wheeled UGVs, they are able to converge to certain configuration with velocity approaching zero. In this case, the problem formulation is reduced to $\mathcal{L}^{+}(\bm{g}_{0})=\{\bm{g}_{d}\}$ and $\lim_{t\to\infty}\bm{g}(t)=\bm{g}_{d}$. Besides, there exist mechanical constraints and task requirements that require non-zero velocity at the target configuration, such as fixed-wing UAVs and glider AUVs on target tracking missions. It makes sense that these robots are guided to the target positive limit set that contains the target configuration, i.e., $\left\{\bm{g}_{d}\right\}\subset\mathcal{L}_{d}^{+}(\bm{g}_{0})$ and hence they achieve the target configuration repeatedly, that is, $\exists\{t_{k}\}_{k=1}^{\infty}\to\infty$ such that $\bm{g}(t_{k})=\bm{g}_{d}$ as $k\to\infty.$

### v2

@@ Abstract
Vector fields are advantageous in handling nonholonomic motion planning, as they provide the robot with reference orientation across the workspace. However, additionally incorporating curvature constraints presents challenges due to the interconnection between the design of the curvature-bounded vector field and the tracking controller under limited actuation. In this paper, we present a novel framework to co-develop the vector field and the control laws, guiding the nonholonomic robot to the target configuration with curvature-bounded trajectory. First, we formulate the problem by introducing the target positive limit set, which allows the robot to either converge to or pass through the target configuration, depending on its dynamics and the specific tasks. Next, we construct a curvature-constrained vector field (CVF) via blending and embedding elementary flows in the workspace. To track such CVF, the saturated control laws with a dynamic gain are proposed, under which the tracking error’s magnitude decreases even when saturation occurs. Under the control laws, kinematically constrained nonholonomic robots are guaranteed to track the reference CVF and converge to the target positive limit set with bounded trajectory curvature. Numerical simulations show that the proposed CVF method outperforms other vector-field-based algorithms. Experiments on Ackermann UGVs and semi-physical fixed-wing UAVs demonstrate that the method can be effectively implemented in real-world scenarios.
@@ I Introduction
Motion planning, a fundamental problem in robotics, is critical for fulfilling various robotic tasks, such as autonomous driving, aerial surveillance, and underwater exploration. A key objective in motion planning is to provide feasible reference trajectories and control inputs to guide the robot to the target configuration. To achieve feasibility and efficiency, motion planning algorithms should take into account the robot’s kinematic constraints, among which the nonholonomic and curvature constraints are frequently encountered. Nonholonomic constraints are typically imposed by the lateral velocity of robots being zero [egeland1996feedback, li2013finite, he2024roto], under which the longitudinal linear velocity and angular velocity are considered as the control inputs in motion planning [panagou2014motion, levin2019real, aiello2022fixed]. In addition, as a result of the limited actuation of nonholonomic robots, such as unmanned ground vehicles (UGVs) with minimum turning radius [balluchi1996optimal, chen2019shortest], unmanned aerial vehicles (UAVs) with maximum lateral acceleration [shima2002time, taub2013intercept], and autonomous underwater vehicles (AUVs) with limited control moment [du2016robust, li2019robust], the trajectory curvature of these robots is subject to an upper bound, termed the curvature constraint. The coexistence of the nonholonomic and curvature constraints poses significant challenges for motion planning, due to the tightly coupled position and orientation kinematics as well as the robots’ limited actuation.
@@
Existing motion planning methods addressing the above constraints fall into several categories, each of which has limitations that hinder its practical implementation. Search-based methods are widely used, but they require post-processing and high computational cost to handle kinematic constraints [murray1993nonholonomic, karaman2011sampling, song2018t]. Optimization-based methods can effectively manage various constraints, but often suffer from issues related to local minima and substantial computational burdens [Hussein2008Optimal, duan2014planning, xu2021autonomous]. Although the Dubins curve-based method is a well-studied algorithm for generating curvature-bounded trajectories for nonholonomic robots, it does not incorporate control input design for path following [dubins1957curves, chen2020dubins, chen2023elongation]. Moreover, the common drawback of the aforementioned methods is that they operate in an open-loop manner, necessitating replanning when deviations occur.
@@
Unlike the preceding open-loop methods, the vector field (VF)-based approaches specify the desired orientation based on the nonholonomic robot’s position by assigning a vector to each point in the workspace. One classic way to generate VFs is by calculating the gradient of potential functions, known as the artificial potential field (APF) [tanner2003nonholonomic, pathak2005integrated, masoud2009harmonic]. However, this method is prone to local minima, which possibly leads to deadlock. To address this problem, Lindemann et al. utilize cell decomposition approach to design a non-gradient VF for each polygonal cell and achieve convergence to the target position [lindemann2006real, lindemann2007smooth, lindemann2009simple]. With dipole-like VF for guidance, Panagou et al. further derive the tracking control laws for the VF to stabilize a nonholonomic robot [panagou2010dipole], and coordinate the motion of multiple unicycles [Panagou2017Distributed]. He et al. [he2025novel] generalize dipole-like VF to 3D motion planning, guiding multiple nonholonomic robots to their target positions with a specific heading direction except for a singular set of zero measures. To address the issues that arise from the singularity of VFs, singularity-free VFs and control laws for fixed-wing UAVs are proposed to achieve path following and motion coordination in [yao2021singularity, yao2023guiding].
Although the above VFs naturally address the nonholonomic constraint, few existing works incorporate the curvature constraint into the VF-based motion planning. The challenges in addressing both the nonholonomic and curvature constraints lie in at least two aspects. On the one hand, designing a curvature-constrained vector field (CVF) is far from trivial, particularly for the simultaneous position and orientation planning problem as discussed in [he2024simultaneous].
When the robot’s orientation aligns with the VF, the robot’s trajectory is the same as the VF’s integral curves by moving along the VF. To ensure the boundedness of the curvature of the trajectory and the continuity of the control inputs, the curvature of the VF should be not only bounded but also continuous[scheuer1997continuous]. Furthermore, navigating towards the target configuration requires the VF to coincide with the target orientation while approaching the target position, which becomes challenging since the adjustment of the orientation of the VF is limited by the curvature constraint. For example, the VFs proposed in [lau2015fluid] would converge to the desired position but do not guarantee alignment with a specified orientation, further highlighting the difficulty of VF design in simultaneous position and orientation planning.
On the other hand, the design of tracking control laws for the CVF becomes significantly more difficult under the nonholonomic and curvature constraints. In most cases, a nonholonomic robot’s orientation differs from the CVF, leading to the deviation of its trajectory from the CVF’s integral curve. Moreover, the curvature constraint imposes an upper bound proportional to the linear velocity on the angular velocity, restricting the robot from aligning its orientation with the CVF and hence exacerbating the deviation. Take [pothen2017curvature, shivam2021curvature], for example, although their VFs have prescribed curvature limits, the control laws designed to track these VFs do not account for the limited actuation mentioned above. Hence, the actual trajectories may violate the curvature constraint when deviations occur. To ensure the existence of feasible control laws that can track the VF under curvature constraint, the co-design of the VF and control laws is crucial. Due to the aforementioned challenges, developing a feedback motion planning method to guide a nonholonomic robot with curvature constraints to the target configuration remains unsolved.
@@
In this paper, we formulate the simultaneous position and orientation planning as guiding the curvature-constrained nonholonomic robot to the target positive limit set, a positive limit set containing the target configuration. To achieve this objective, we aim to find the reference trajectories and control inputs that satisfy the curvature constraint while converging to the target positive limit set. This formulation is more general and applies to real-world scenarios, including Ackermann UGV converging to a configuration, and fixed-wing UAV passing through a configuration with nonzero velocity. In order to address the difficulties mentioned above, we propose a unified VF-based motion planning framework, under which the VF provides the curvature-bounded reference trajectories and the control laws ensure that the nonholonomic robot tracks the proposed VF under the curvature constraint.
The theoretical contributions of this paper are summarized as follows:
@@
  • We construct a curvature-constrained vector field (CVF) with a stable limit cycle as the target positive limit set. A novel VF is proposed by partitioning the workspace and embedding the blended source, sink and vortex flows into different regions with conveniently tunable parameters. Moreover, we derive the sufficient condition ensuring that the CVF’s integral curves have a prescribed curvature bound and serve as the reference trajectories converging to the limit cycle except at an isolated singular point. Compared with the dipole-like VF in [panagou2010dipole, he2025novel] with singular target position and the curvature-constrained VF in [shivam2021curvature, pothen2017curvature], this construction strategy reduces the effect of singularity during trajectory convergence, while allowing a flexible selection of CVF parameters, facilitating the co-design of CVF and control laws.
@@
  • We design novel control laws comprising the linear velocity with specified bounds adaptive to the planning objective and the saturated angular velocity, which ensures that the actual trajectory curvature is also bounded when tracking the CVF. To guarantee that the closed-loop system tracks the CVF under saturation of angular velocity, we propose a state-dependent dynamic gain that responds to changes in the CVF. In this way, the orientation error between the robot and CVF is monotonically stabilized even when saturation occurs.
@@
To validate the theoretical results and facilitate reproducibility, we conduct and open source the implementations, including the numerical simulations and hardware experiments [^1] Implementation details and source code are available at https://github.com/Y1kee/CVF-for-Nonholonomic-Motion-Planning, as outlined below.
  • We carry out numerical simulations demonstrating supportive phenomena for the theoretical propositions and Monte Carlo experiments comparing our method to other VF-based algorithms in [Panagou2017Distributed, he2024simultaneous, pothen2017curvature, yao2021singularity]. The presented approach outperforms the existing methods by achieving a 100% success rate in ensuring convergence under curvature constraints.
  • We implement the proposed planning algorithm on the Ackermann UGV featuring a minimal turning radius. The experiment results show that they successfully converge to different destinations from various initial conditions, demonstrating that the proposed feedback motion planning algorithm is robust against the dynamics uncertainties.
  • We extend the proposed method to 3D fixed-wing UAVs with coordinated turn constraint by deriving the attitude and thrust control laws. In the hardware-in-the-loop (HIL) experiments with target configuration switching during the flight, the UAV reaches its destination under the extended control laws. The HIL results highlight that our algorithm given by closed-form expressions enables real-time motion planning despite limited onboard computational resources.
The paper proceeds as follows. Section II presents the nonholonomic kinematics with curvature constraint and formulates the motion planning problem. Then, Section III shows how to build the CVF and demonstrates its desired properties of convergence and bounded curvature. To track the proposed CVF, saturated control laws with dynamic gains are introduced in Section IV, where almost global convergence under curvature constraint is established. Section V conducts numerical simulations to validate the theoretical results and compares our algorithm with other existing VF-based motion planners. In Section VI, the proposed method is further evaluated through experiments using an Ackermann UGV and a semi-physical fixed-wing UAV with curvature-bounded trajectories. Finally, Section VII concludes this paper.
@@ II-A Modeling of Kinematics and Constraints
In this paper, we let $\bm{g}=(\bm{p},\theta)$ denote the configuration of a 2D mobile robot, where $\bm{p}=[x,y]^{\intercal}\in\mathbb{R}^{2}$ is the position and $\theta\in[0,2\pi)$ is the orientation. For the considered configuration, the nonholonomic constraint is imposed by confining the translational motion along the robot’s heading direction, i.e., the $x$-axis of the body-fixed frame. In this case, the most commonly used nonholonomic kinematics are expressed as
@@
where the control inputs consist of the linear velocity $v_{x}$ and the angular velocity $\omega$.
Due to the limitation on the steering mechanism of nonholonomic robots, the trajectory curvature is subject to an upper bound, known as the curvature constraint. For kinematics depicted by (1), the curvature constraint is written as
@@
where $\kappa$ is the curvature of robot’s actual trajectory and $\bar{\kappa}$ is the maximum curvature. Additionally, the minimum turning radius of the nonholonomic robot is denoted by
@@
The above kinematic model captures the essential constraints for us to conduct rigorous analysis of the algorithm to be proposed, and serves as a foundation extendable to more complex robotic systems. The experiment results demonstrate that the proposed method is applicable to UGVs with Ackermann steering geometry and fixed-wing UAVs with banking turn control featuring cruising speed constraints.
@@ II-B Problem Formulation
In real-world scenarios, mobile robots are usually required to reach the target position with predefined orientation to accomplish specific tasks. For robots with positive lower bound on linear velocity due to task requirements or dynamics constraints, the convergence to a specific configuration is not always achievable. For a more general problem formulation, we employ the concept of positive limit set in [khalil2002nonlinear, Ch. 4.2] and define the target positive limit set as follows.
@@
According to the above definitions, a controlled robot could approach the target configuration by converging to the target positive limit set. To achieve this objective, the motion planner aims to derive curvature-bounded reference trajectories that converge to the target positive limit set, along with control inputs that track these reference trajectories. In this way, the nonholonomic system (1) subject to the curvature constraint (2) converges to the target positive limit set, i.e.,
@@
For nonholonomic robots like wheeled UGVs, they are able to converge to certain configuration with velocity approaching zero. In this case, the problem formulation is reduced to $\mathcal{L}^{+}(\bm{g}_{0})=\{\bm{g}_{d}\}$ and $\lim_{t\to\infty}\bm{g}(t)=\bm{g}_{d}$. Besides, there exist dynamics constraints and task requirements that require non-zero velocity at the target configuration, such as fixed-wing UAVs and glider AUVs. It makes sense to guide these robots to the target positive limit set that contains the target configuration, i.e., $\left\{\bm{g}_{d}\right\}\subset\mathcal{L}_{d}^{+}(\bm{g}_{0})$, and hence they reach the target configuration repeatedly, i.e., $\exists\{t_{k}\}_{k=1}^{\infty}\to\infty$ such that $\bm{g}(t_{k})=\bm{g}_{d}$ as $k\to\infty.$